Differentiate the following with respect to x

e^x/log x
1
Expert's answer
2020-09-28T20:55:46-0400

"let \space y= \frac{e^x}{logx}\\differentiating \space both \space sides \space w.r.t \space to \space 'x'...we \space get"

"\frac{dy}{dx}= \frac{logx*\frac {d}{dx}e^x - e^x*\frac{d}{dx}logx}{(logx)^2}\\" ( using quotient rule of differentiation i.e "\frac {d}{dx}(u/v) = \frac{v*\frac{du}{dx} - u*\frac{dv}{dx}
}{v^2}" )


"or, \frac{dy}{dx}= \frac{logx*e^x - e^x*1/x}{(logx)^2}\\"


"or, \frac{dy}{dx}=\frac{ e^x [logx \space - \space 1/x]}{(logx)^2}" (ANSWER)